OpenStudy (anonymous):

Medal&Fan There are two consecutive even numbers such that five times the smaller minus three times the greater is 6. What is the smaller number? Select one: a. 2 b. 4 c. 6 d. 8

OpenStudy (twopointinfinity):

the answer is C: 6

OpenStudy (twopointinfinity):

If x is an even number, then x+2 is the next even number (because if you only add 1 to an even number then you have an odd number). X+2 is the greater number. five times the smaller number can be expressed as 5x. Three times the greater number is expressed as 3(x+2). The equation is then: 5x - 3(x+2) = 6 using the distributive law the equation becomes: 5x -3x -6 = 6 combine like terms: 2x - 6 = 6 then: 2x = 12 and x = 6 x+2 = 8 5(6) - 3(8) = 30 - 24 = 6

OpenStudy (anonymous):

wow. thankss.c: